OpenStudy (anonymous):

Put the equation in the form ax^2+bx+c=0 7-3x^2= (x+2)^2-6

OpenStudy (anonymous):

My I have permission to help?

OpenStudy (anonymous):

lol hero

OpenStudy (anonymous):

I'm not trying to be funny

OpenStudy (anonymous):

sureeeeee

OpenStudy (anonymous):

1. Expand (x+2)^2: 7 - 3x^2 = x^2 +4x + 4 -6 2. Add 3x^2 to both sides: 7 = 4x^2 + 4x -2 3. Subtract 7 from both sides: 0 = 4x^2 +4x -9

OpenStudy (anonymous):

You see....I'm only here to help

OpenStudy (anonymous):

That's good. All I want is help.

OpenStudy (anonymous):

wow